complete order pair for 2x+y=10
(5, )( ,10)( ,-2)(0, )

Answer by elima(1433)

complete order pair for 2x+y=10
(5, )( ,10)( ,-2)(0, )
===================================
2(5)+y=10
10+y=10
y=0
(5,0)
--------------
2x+10=10
2x=0
x=0
(0,10)
----------------
2x+-2=10
2x=12
x=6
(6,-2)
----------------
2(0)+y=10
o+y=10
y=10
(0,10)
:)